# Galactic bars are already mature at Cosmic Noon: bar strength and flatness at $z\sim 1.5$

### v1

@@ 6 Summary
In this work, we explore the nature of bars in a sample of nine massive galaxies ($M_{\mathrm{*}}/M_{\mathrm{\odot}}=10^{10.6-11.4}$) at $z_{\mathrm{spec}}\sim 1.5$. We use rest-frame near-IR JWST/NIRCam F [^4] W imaging from the COSMOS-Web survey (casey23). Leveraging the high spatial resolution, we perform a suite of morphological analyses to derive a range of bar properties:
@@
  • We find that five of the nine galaxies host flat, mature bars with $n<0.4$. Of the remaining four, only two bars have exponential profiles ($n>0.8$), indicating young bars. Hence, the bars at $z\sim 1.5$ have survived long enough ($\gtrsim 1-2\,\rm Gyr$, based on N-body simulations) to develop flat profiles (Sec. 5.1).
@@
  • For the bars with measured $Q_{\mathrm{b}}$, we find a marginal decrease of the value with the bar Sérsic index (Sec. 5.2), consistent with the expectation that bars grow stronger as they mature.
@@
  • Through comparison with simulations, we conclude that the mature bars at $z\sim 1.5$ suggest the presence of settled, dominant disks by $z\sim 2.5-3.5$ (Sec. 5.3).
@@
The authors are very grateful for the valuable comments by the anonymous referee. BSK was supported by the China Postdoctoral Science Foundation under Grant Number 2025M773186. LCH was supported by the National Science Foundation of China (12233001) and the China Manned Space Program (CMS-CSST-2025-A09). This work was made possible by utilizing the CANDIDE cluster at the Institut d’Astrophysique de Paris, which was funded through grants from the PNCG, CNES, DIM-ACAV, and the Cosmic Dawn Center and maintained by Stephane Rouberol. BSK would like to thank Jinyi Shangguan for the valuable discussions. Finally, BSK thanks Xueling Wang for the constant support, without which this work would not have been possible. {contribution}BSK was responsible for most of the aspects of this manuscript. LCH and FB provided key insights into the interpretation of the results. JDS, ED provided valuable comments during the compilation of the results. XD and SY provided help with components of the code that was used for the analysis. Finally, all other authors provided valuable comments during the preparation of the manuscript.

### v2

@@ VI Summary
In this work, we explore the nature of bars in a sample of nine massive galaxies ($M_{\mathrm{*}}/M_{\mathrm{\odot}}=10^{10.6-11.4}$) at $z_{\mathrm{spec}}\sim 1.5$. We use rest-frame near-IR JWST/NIRCam F [^4] W imaging from the COSMOS-Web survey (C. M. Casey et al., 2023). Leveraging the high spatial resolution, we perform a suite of morphological analyses to derive a range of bar properties:
@@
  • We find that five of the nine galaxies host flat, mature bars with $n<0.4$. Of the remaining four, only two bars have exponential profiles ($n>0.8$), indicating young bars. Hence, the bars at $z\sim 1.5$ have survived long enough ($\gtrsim 1-2\,\rm Gyr$, based on N-body simulations) to develop flat profiles (Sec. V.1).
@@
  • For the bars with measured $Q_{\mathrm{b}}$, we find a marginal decrease of the value with the bar Sérsic index (Sec. V.2), consistent with the expectation that bars grow stronger as they mature.
@@
  • Through comparison with simulations, we conclude that the mature bars at $z\sim 1.5$ suggest the presence of settled, dominant disks by $z\sim 2.5-3.5$ (Sec. V.3).
@@
The authors are very grateful for the valuable comments by the anonymous referee. BSK was supported by the China Postdoctoral Science Foundation under Grant Number 2025M773186. LCH was supported by the National Science Foundation of China (12233001) and the China Manned Space Program (CMS-CSST-2025-A09). This work was made possible by utilizing the CANDIDE cluster at the Institut d’Astrophysique de Paris, which was funded through grants from the PNCG, CNES, DIM-ACAV, and the Cosmic Dawn Center and maintained by Stephane Rouberol. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1727. The specific observations analyzed in this work can be accessed via DOI: 10.17909/p4e8-bz21. BSK would like to thank Jinyi Shangguan for the valuable discussions. Finally, BSK thanks Xueling Wang for the constant support, without which this work would not have been possible.
{contribution}
BSK was responsible for most of the aspects ofJDS, ED,MD and AP provided valuable comments during the compilation of the results this manuscript. LCH and FB provided key insights into the interpretation of the results. . XD and SY provided help with components of the code that was used for the analysis. Finally, all other authors provided valuable comments during the preparation of the manuscript.